# Planar percolation and the loop $\mathrm{O}(n)$ model

### v2

@@ Abstract.
We show that a large class of site percolation processes on any planar graph contains either zero or infinitely many infinite connected components. The assumptions that we require are: tail triviality, positive association (FKG) and that the set of open vertices is stochastically dominated by the set of closed ones. This covers the case of Bernoulli site percolation at parameter $p\leq 1/2$ and resolves Conjecture $8$ from the work of Benjamini and Schramm from 1996. Our result also implies that $p_{c}\geq 1/2$ for any invariantly amenable unimodular random rooted planar graph.
@@ 1. Introduction
Modern statistical physics has its origin in the study of the Lenz-Ising model and its phase transition. This model of ferromagnetic interactions was first defined more than a century ago. It was later generalized to a large class of spin lattice models meant as a discrete approximation to Euclidean geometry. A few decades later, Broadbent-Hammersley [BH57] introduced Bernoulli percolation. This is a model of random subgraphs of a given (crystallographic) lattice undergoing a phase transition in terms of the existence of an infinite connected component. It was then discovered that percolation models provide a graphical representation for correlations in Ising-like spin models. Since then, percolation models have been at the center of the mathematical approach to phase transitions; see [Gri99] for a classical manuscript and [Man25] for a recent survey.
Special attention has been lavished upon two-dimensional lattice models, where the scaling limits of lattice approximations are expected to be described by a conformal field theory. These emergent symmetries have lead to the wide-open conjecture that certain interfaces converge to the Schramm–Loenwer Evolution (SLE). This has been proved in only a handful of models, including critical Ising model on isoradial graphs [Smi10, CS11] and critical site percolation on the triangular lattice [Smi01b, Smi01a].
Much of the study of percolation is focused on vertex (quasi-)transitive graphs, where ergodic-theoretic arguments can be used to restrict the possible behavior of Bernoulli percolation. Indeed, denote by $N_{\infty}$ the number of infinite connected components. Then, it is straightforward to show that, on any connected transitive graph, $N_{\infty}$ is almost surely constant, and can only be zero, one, or infinity. If the graph is amenable, a powerful argument of Burton and Keane [BK89] rules out the case $N_{\infty}=\infty$. Also, if $N_{\infty}=1$ for Bernoulli percolation of some parameter $p$, then the same holds for any $q>p$ [HP99]. Both statements are false if one considers general planar graphs; see Fig. 1 for examples where $N_{\infty}$ has a non-trivial distribution and the property $N_{\infty}=1$ is non-monotone in $p$.
@@
In this work we study general locally finite planar graphs and rule out the possibility that $1\leq N_{\infty}<\infty$ for Bernoulli site percolation whenever $p\leq 1/2$. In fact, we prove this result for a more general class of site percolation models (Theorem 1). We emphasize that Theorem 1 does not require any symmetry and holds true also when planar embeddings of $G$ have any number of accumulation points. The case of Bernoulli percolation at $p=1/2$ settles Conjecture $8$ in the seminal work of Benjamini and Schramm [BS96]. When $G$ is quasi-transitive, this conjecture has been established recently by Grimmett and Li [GL25b].
In instances where one can rule out the scenario where $N_{\infty}=\infty$ a priori, Theorem 1 implies the absence of infinite connected components and provides the optimal lower bound on the percolation threshold $p_{c}$. We show that $p_{c}\geq 1/2$ for any invariantly amenable unimodular random rooted planar graph (Corollary 1.1). This improves on a uniform lower bound proved by Peled [Pel20] and can be thought of as a generalization of the no-coexistence theorems of Zhang (unpublished, described in [Gri99, Lemma 9.12]) and Sheffield [She05] (see also [DRT19]). Such results have been very useful in proving delocalization of height functions [CPST21, Lam21, GM21] which allowed for new proofs of the continuity of the phase transition [GL25a] and the existence of a Berezinskii–Kosterlitz–Thouless (BKT) phase [EL23, AHPS21]; see [FS81, DST17] for the original proofs. Since we do not rely on the symmetries of the graphs, our method opens a path to extend all these results to more general settings.
Corollary 1.2 extends the statement $p_{c}\geq 1/2$ of Corollary 1.1 to divide and color models, a general class which includes the Ising and fuzzy Potts models as particular cases. As we show below, the loop $\mathrm{O}(n)$ model also fits in this class. This model is supported on collections of non-intersecting cycles (loops) on the hexagonal lattice and has parameters $n>0$ (loop-weight) and $x>0$ (edge-weight). It was introduced in 1981 [DMNS81] and is difficult to study due to the absence of monotonicity. This model is conjectured to undergo a phase transition in terms of loop lengths for all $n\in(0,2]$ [Nie82]: macroscopic loops when $x\geq x_{c}(n)$ versus exponential decay when $x<x_{c}(n)$.
We use Theorem 1 via Corollary 1.2 to establish a large portion of the conjectured phase diagram of the loop $\mathrm{O}(n)$ model (Theorem 2). Specifically, we prove that each face is surrounded by infinitely many loops when $n\in[1,2],x\in[1/\sqrt{2},1]$. This covers nearly all previously known results of this sort [DGPS21, GM21, CGHP25, GL25a]. Together with the dichotomy shown in [DGPS21], our result implies the macroscopic behavior when $n\in[1,2],x\in[1/\sqrt{2},1/\sqrt{n}]$.
@@ 1.1. Site percolation models on planar graphs
In the last item, the random variable $1-\sigma$ is defined by $(1-\sigma)_{v}:=1-\sigma_{v}$ for each $v\in V$. By the Kolmogorov zero-one law and the Harris inequality [Har60], Bernoulli site percolation satisfies the first two properties; whenever $p\leq 1/2$, it also satisfies the third.
@@
The study of percolation on general planar graphs was initiated by Benjamini and Schramm [BS96]. That celebrated work asks which properties of Bernoulli site percolation are particular to two-dimensional lattices and which ones hold in general. Their Conjecture $8$ states the claim of Theorem 1 for Bernoulli site percolation at $p=1/2$. As mentioned above, this model satisfies the assumptions of Theorem 1 and, thus, we settle the conjecture completely. The statement at $p=1/2$ for quasi-transitive planar graphs is the content of [GL25b, Theorem 1.9].
Theorem 1 is also linked to [BS96, Conjecture $7$] that states that, for any planar graph of minimal degree at least $7$, $p_{c}<1/2$ and $p_{u}\geq 1-p_{c}$, where $p_{c}$ and $p_{u}$ are defined by
@@
In the case where $G$ can be embedded in the plane with no accumulation points, this conjecture has been confirmed by Haslegrave and Panagiotis [HP21] ($p_{c}<1/2$) and Li [Li23] ($p_{u}\geq 1-p_{c}$). We expect that $p_{u}\geq 1-p_{c}$ holds also without any assumptions on the degrees of the graph. Indeed, our result implies that
$p_{u}\geq 1/2$
for any infinite locally finite planar graph, and $p_{u}=p_{c}\geq 1/2$ if $G$ is planar, unimodular, and invariantly amenable (Corollary 1.1). It is classical [BS01] that $p_{u}\geq 1-p_{c}$ if $G$ is planar, unimodular, and invariantly nonamenable.
Finally, we want to emphasize two robust features of the arguments presented in this work. First, our methods are not restricted to Bernoulli site percolation. Instead, we rely on the FKG inequality and tail triviality. This extends the study of planar percolations to this more general class of processes that may have additional applications. Second, our methods do not use any special properties of a ‘nice’ choice of embedding of the planar graph $G$, such as a circle packing. Indeed, unlike most earlier works on planar percolations, we do not require the existence of a proper planar embedding (i.e. without accumulation points). We hope that our approach can be used to remove this extraneous assumption also in other cases.
@@ 1.2. Corollaries for percolation on unimodular invariantly amenable graphs
Classical results of Burton and Keane [BK89] and Zhang (see [Gri99, Chapter 11]) imply that $p_{c}=p_{u}\geq 1/2$ on any (quasi-)transitive amenable graph. Using Theorem 1, this can be extended to the case of unimodular invariantly amenable graphs (Corollary 1.1). Let $\mathcal{G}_{\bullet}$ (resp. $\mathcal{G}_{\bullet\bullet}$) be the set of all isomorphism classes of locally finite planar graphs together with a distinguished vertex (resp. two distinguished vertices) equipped with the local topology. A rooted random planar graph $(G,\rho)$ is a random variable with values in $\mathcal{G}_{\bullet}$. We say that $(G,\rho)$ is unimodular if, for any Borel function $f\colon\mathcal{G}_{\bullet\bullet}\to[0,\infty]$, we have
@@
A bond percolation on a graph $G=(V,E)$ is a random variable $\omega$ that takes values in $\{0,1\}^{E}$. We identify $\omega$ with the spanning subgraph of $G$ given by $\{\omega=1\}$; we say that it is finitary if all connected components are finite, almost surely, and invariant if $(G,\rho,\omega)$ is unimodular in a sense that (MTP) holds also when $f$ depends on both $(G,\rho)$ and $\omega$ (see [AHNR18, Section 2] for more details). For $v\in V$, define $K_{\omega}(v)$ as the connected component of $v$ in $\omega$; for $U\subseteq V$, define $\partial U$ as the set of vertices in $U$ that are adjacent to $V\smallsetminus U$. We say that $(G,\rho)$ is (vertex) invariantly amenable if
@@
Our definition via inner vertex boundaries agrees with [AL07] and is different from [AHNR18], which uses edge boundaries. Note that if a graph is edge invariantly amenable in the sense of [AHNR18], then it is also vertex invariantly amenable. Moreover, when the root has a finite expected degree, being unimodular and (edge) invariantly amenable is equivalent to being the Benjamini–Schramm (or local) limit of finite planar maps [LT80, AHNR18].
@@
Previously, Peled [Pel20] proved the existence of some $c>0$ such that $p_{c}>c$ for the Benjamini–Schramm limit of arbitrary finite planar maps; that work also shows that the circle packing of such graphs exhibits exponential decay at all $p<c$. We do not know whether all such graph are (vertex) invariantly amenable, but, as mentioned above, the implication is true when the root has a finite expected degree. Corollary 1.1 thus gives that $p_{c}(G)\geq 1/2$ if $(G,\rho)$ is the Benjamini–Schramm limit of finite planar maps and $\rho$ has a finite expected degree. This bound is optimal since the site percolation on the triangular lattice has $p_{c}=1/2$ [Kes80].
We further extend the statement to the framework of divide and color models (Corollary 1.2). For a graph $G=(V,E)$, we define $\mathrm{ER}(V)$ as the set of all equivalence relations on $V$. Given $\mu$ a probability measure on $\mathrm{ER}(V)$, a divide and color model of parameter $p\in[0,1]$ is a site percolation $\sigma$ on $G$ defined as follows. First, one samples a partition $P=\{P_{i}\}_{i=1}^{\infty}$ from $\mu$. Then, for each $i\in\mathbb{N}$, we assign all vertices of $P_{i}$ the value $\sigma=1$ with probability $p$ and $\sigma=0$ otherwise, independently of all other classes. The model was first introduced by Häggström [Häg01] in the case when $\mu$ is Bernoulli bond percolation; the critical value function in this case was studied in [BBT13b, BBT13a]. The Gibbs properties of the measure when $\mu$ is the random-cluster model was considered in [Bál10].
A version with a general partition $\mu$ and several colors is due to Steif and Tykesson [ST19]. Here we focus on a two coloring, which recovers the Ising, fuzzy Ising, and Voter models as particular cases and allows us to apply Theorem 1 to the loop $\mathrm{O}(n)$ model (Theorem 2). We call a random partition $P$ of $V$ finitary if its elements are finite, almost surely, and invariant if $(G,\rho,P)$ is unimodular.
@@
Corollaries 1.1 and 1.2 follow from Theorem 1 once one rules out infinitely many infinite components. The latter statement was proven in a number of cases: the classical argument of Burton and Keane [BK89] treats shift invariant site percolation processes on $\mathbb{Z}^{d}$ with the finite energy property. This argument was then extended in [AL07, Sections $6,8$] based on [LS99, BLPS99, BLPS01] (see also [AHNR18, Theorem 5.9]) to insertion-tolerant invariant bond percolation processes on unimodular graphs. In [LP17, Theorem 7.9], insertion-tolerant invariant site and bond percolation processes on transitive graphs were treated. The arguments in these works use insertion-tolerance only to construct an infinite component with at least three ends, so the assumption can be weakened. The combination of these arguments extends the statement to the setting of Corollaries 1.1 and 1.2; we provide short proofs for completeness.

### v3

@@ Abstract.
We show that a large class of site percolation processes on any planar graph contains either zero or infinitely many infinite connected components. The assumptions that we require are: tail triviality, positive association (FKG), and that the set of open vertices is stochastically dominated by the set of closed ones. This covers the case of Bernoulli site percolation at parameter $p\leq 1/2$ and resolves Conjecture $8$ from the work of Benjamini and Schramm from 1996. Our result also implies that $p_{c}\geq 1/2$ for any invariantly amenable unimodular random rooted planar graph.
@@ 1. Introduction
Modern statistical physics has its origin in the study of the Lenz-Ising model and its phase transition. This model of ferromagnetic interactions was first defined more than a century ago. It was later generalized to a large class of spin lattice models meant as a discrete approximation to Euclidean geometry. A few decades later, Broadbent-Hammersley [11] introduced Bernoulli percolation. This is a model of random subgraphs of a given (crystallographic) lattice undergoing a phase transition in terms of the existence of an infinite connected component. It was then discovered that percolation models provide a graphical representation for correlations in Ising-like spin models. Since then, percolation models have been at the center of the mathematical approach to phase transitions; see [30] for a classical manuscript and [48] for a recent survey.
Special attention has been lavished upon two-dimensional lattice models, where the scaling limits of lattice approximations are expected to be described by a conformal field theory. These emergent symmetries have lead to the wide-open conjecture that certain interfaces converge to the Schramm–Loenwer Evolution (SLE). This has been proved in only a handful of models, including critical Ising model on isoradial graphs [55, 17] and critical site percolation on the triangular lattice [53, 54].
Much of the study of percolation is focused on vertex (quasi-)transitive graphs, where ergodic-theoretic arguments can be used to restrict the possible behavior of Bernoulli percolation. Indeed, denote by $N_{\infty}$ the number of infinite connected components. Then, it is straightforward to show that, on any connected transitive graph, $N_{\infty}$ is almost surely constant, and can only be zero, one, or infinity. If the graph is amenable, a powerful argument of Burton and Keane [12] rules out the case $N_{\infty}=\infty$. Also, if $N_{\infty}=1$ for Bernoulli percolation of some parameter $p$, then the same holds for any $q>p$ [32]. Both statements are false if one considers general planar graphs; see Fig. 1 for examples where $N_{\infty}$ has a non-trivial distribution and the property $N_{\infty}=1$ is non-monotone in $p$.
@@
In this work we study general locally finite planar graphs and rule out the possibility that $1\leq N_{\infty}<\infty$ for Bernoulli site percolation whenever $p\leq 1/2$. In fact, we prove this result for a more general class of site percolation models (Theorem 1). We emphasize that Theorem 1 does not require any symmetry, and also holds true when planar embeddings of $G$ have any number of accumulation points. The case of Bernoulli percolation at $p=1/2$ settles Conjecture $8$ in the seminal work of Benjamini and Schramm [8]. When $G$ is quasi-transitive, this conjecture has been established recently by Grimmett and Li [31].
In instances where one can rule out the scenario where $N_{\infty}=\infty$ a priori, Theorem 1 implies the absence of infinite connected components and provides the optimal lower bound on the percolation threshold $p_{c}$. We show that $p_{c}\geq 1/2$ for any invariantly amenable unimodular random rooted planar graph (Corollary 1.1). This improves on a uniform lower bound proved by Peled [51] and can be thought of as a generalization of the no-coexistence theorems of Zhang (unpublished, described in [30, Lemma 9.12]) and Sheffield [52] (see also [21]). Such results have been very useful in proving delocalization of height functions [14, 42, 28] which allowed for new proofs of the continuity of the phase transition [27] and the existence of a Berezinskii–Kosterlitz–Thouless (BKT) phase [24, 1]; see [26, 22] for the original proofs. Since we do not rely on the symmetries of the graphs, our method opens a path to extend all these results to more general settings.
Corollary 1.2 extends the statement $p_{c}\geq 1/2$ of Corollary 1.1 to divide and color models, a general class which includes the Ising and fuzzy Potts models as particular cases. As we show below, the loop $\mathrm{O}(n)$ model also fits in this class. This model is supported on collections of non-intersecting cycles (loops) on the hexagonal lattice and has parameters $n>0$ (loop-weight) and $x>0$ (edge-weight). It was introduced in 1981 [19] and is difficult to study due to the absence of monotonicity. This model is conjectured to undergo a phase transition in terms of loop lengths for all $n\in(0,2]$ [49]: macroscopic loops when $x\geq x_{c}(n)$ versus exponential decay when $x<x_{c}(n)$.
We use Theorem 1 via Corollary 1.2 to establish a large portion of the conjectured phase diagram of the loop $\mathrm{O}(n)$ model (Theorem 2). Specifically, we prove that each face is surrounded by infinitely many loops when $n\in[1,2],x\in[1/\sqrt{2},1]$. This covers nearly all previously known results of this sort [20, 28, 18, 27]. Together with the dichotomy shown in [20], our result implies the macroscopic behavior when $n\in[1,2],x\in[1/\sqrt{2},1/\sqrt{n}]$.
@@ 1.1. Site percolation models on planar graphs
In the last item, the random variable $1-\sigma$ is defined by $(1-\sigma)_{v}:=1-\sigma_{v}$ for each $v\in V$. By the Kolmogorov zero-one law and the Harris inequality [34], Bernoulli site percolation satisfies the first two properties; whenever $p\leq 1/2$, it also satisfies the third.
@@
The study of percolation on general planar graphs was initiated by Benjamini and Schramm [8]. That celebrated work asks which properties of Bernoulli site percolation are particular to two-dimensional lattices and which ones hold in general. Their Conjecture $8$ states the claim of Theorem 1 for Bernoulli site percolation at $p=1/2$. As mentioned above, this model satisfies the assumptions of Theorem 1 and, thus, we settle the conjecture completely. The statement at $p=1/2$ for quasi-transitive planar graphs is the content of [31, Theorem 1.9].
Theorem 1 is also linked to [8, Conjecture $7$] that states that, for any planar graph of minimal degree at least $7$, $p_{c}<1/2$ and $p_{u}\geq 1-p_{c}$, where $p_{c}$ and $p_{u}$ are defined by
@@
In the case where $G$ can be embedded in the plane with no accumulation points, this conjecture has been confirmed by Haslegrave and Panagiotis [35] ($p_{c}<1/2$) and Li [43] ($p_{u}\geq 1-p_{c}$). Theorem 1 implies that $p_{u}\geq 1/2$ for any infinite locally finite planar graph, and $p_{u}=p_{c}\geq 1/2$ if $G$ is planar, unimodular, and invariantly amenable (Corollary 1.1). It is classical [10] that $p_{u}\geq 1-p_{c}$ if $G$ is planar, unimodular, and invariantly nonamenable. Building on an initial preprint of the current work, Li [44] proved that $p_{u}\geq 1-p_{c}$ for any planar graph with countably many accumulation points in its Freudenthal embedding. She also constructed a planar graph where $p_{u}<1-p_{c}$. This graph must have uncountably many accumulation points in all of its planar embeddings.
Finally, we want to emphasize two robust features of the arguments presented in this work. First, our methods are not restricted to Bernoulli site percolation. Instead, we rely on the FKG inequality and tail triviality. This extends the study of planar percolations to this more general class of processes, which may allow for additional applications. Second, our methods do not use any special properties of a ‘nice’ choice of embedding of the planar graph $G$, such as a circle packing. Indeed, unlike most earlier works on planar percolations, we do not require the existence of a proper planar embedding (i.e. without accumulation points). We hope that our approach can be used to remove this extraneous assumption in other cases as well.
@@ 1.2. Corollaries for percolation on unimodular invariantly amenable graphs
Classical results of Burton and Keane [12] and Zhang (see [30, Chapter 11]) imply that $p_{c}=p_{u}\geq 1/2$ on any (quasi-)transitive amenable graph. Using Theorem 1, this can be extended to the case of unimodular invariantly amenable graphs (Corollary 1.1). Let $\mathcal{G}_{\bullet}$ (resp. $\mathcal{G}_{\bullet\bullet}$) be the set of all isomorphism classes of locally finite planar graphs together with a distinguished vertex (resp. two distinguished vertices) equipped with the local topology. A rooted random planar graph $(G,\rho)$ is a random variable with values in $\mathcal{G}_{\bullet}$. We say that $(G,\rho)$ is unimodular if, for any Borel function $f\colon\mathcal{G}_{\bullet\bullet}\to[0,\infty]$, we have
@@
A bond percolation on a graph $G=(V,E)$ is a random variable $\omega$ that takes values in $\{0,1\}^{E}$. We identify $\omega$ with the spanning subgraph of $G$ given by $\{\omega=1\}$; we say that it is finitary if all connected components are finite, almost surely, and invariant if $(G,\rho,\omega)$ is unimodular in a sense that (MTP) holds also when $f$ depends on both $(G,\rho)$ and $\omega$ (see [3, Section 2] for more details). For $v\in V$, define $K_{\omega}(v)$ as the connected component of $v$ in $\omega$; for $U\subseteq V$, define $\partial U$ as the set of vertices in $U$ that are adjacent to $V\smallsetminus U$. We say that $(G,\rho)$ is (vertex) invariantly amenable if
@@
Our definition via inner vertex boundaries agrees with [2] and is different from [3], which uses edge boundaries. Note that if a graph is edge invariantly amenable in the sense of [3], then it is also vertex invariantly amenable. Moreover, when the root has a finite expected degree, being unimodular and (edge) invariantly amenable is equivalent to being the Benjamini–Schramm (or local) limit of finite planar maps [45, 3].
@@
Previously, Peled [51] showed that there exists $c>0$ such that $p_{c}(G)\geq c$ for any planar, locally finite graph $G$ that may be represented by a circle packing with countably many accumulation points. Thanks to the celebrated work of He–Schramm [36], this class of graphs includes all recurrent simple planar triangulations with countably many ends. If view of [3], the result of [51] also implies $p_{c}(G)\geq c$ for every unimodular (edge) invariantly amenable with finite expected degree, since any circle packing of such graph has at most one accumulation point. We do not know whether all such graph are (vertex) invariantly amenable, but, as mentioned above, the implication is true when the root has a finite expected degree. Corollary 1.1 thus gives that $p_{c}(G)\geq 1/2$ if $(G,\rho)$ is the Benjamini–Schramm limit of finite planar maps and $\rho$ has a finite expected degree. This bound is optimal since the site percolation on the triangular lattice has $p_{c}=1/2$ [39].
We further extend the statement to the framework of divide and color models (Corollary 1.2). For a graph $G=(V,E)$, we define $\mathrm{ER}(V)$ as the set of all equivalence relations on $V$. Given $\mu$ a probability measure on $\mathrm{ER}(V)$, a divide and color model of parameter $p\in[0,1]$ is a site percolation $\sigma$ on $G$ defined as follows. First, one samples a partition $P=\{P_{i}\}_{i=1}^{\infty}$ from $\mu$. Then, for each $i\in\mathbb{N}$, we assign all vertices of $P_{i}$ the value $\sigma=1$ with probability $p$ and $\sigma=0$ otherwise, independently of all other classes. The model was first introduced by Häggström [33] in the case when $\mu$ is Bernoulli bond percolation; the critical value function in this case was studied in [5, 4]. The Gibbs properties of the measure when $\mu$ is the random-cluster model was considered in [6].
A version with a general partition $\mu$ and several colors is due to Steif and Tykesson [56]. Here we focus on a two coloring, which recovers the Ising, fuzzy Ising, and Voter models as particular cases, and allows us to apply Theorem 1 to the loop $\mathrm{O}(n)$ model (Theorem 2). We call a random partition $P$ of $V$ finitary if its elements are finite, almost surely, and invariant if $(G,\rho,P)$ is unimodular.
@@
Corollaries 1.1 and 1.2 follow from Theorem 1 once one rules out infinitely many infinite components. The latter statement was proven in a number of cases: the classical argument of Burton and Keane [12] treats shift invariant site percolation processes on $\mathbb{Z}^{d}$ with the finite energy property. This argument was then extended in [2, Sections $6,8$] based on [47, 7, 9] (see also [3, Theorem 5.9]) to insertion-tolerant invariant bond percolation processes on unimodular graphs. In [46, Theorem 7.9], insertion-tolerant invariant site and bond percolation processes on transitive graphs were treated. The arguments in these works use insertion-tolerance only to construct an infinite component with at least three ends, so the assumption can be weakened. The combination of these arguments extends the statement to the setting of Corollaries 1.1 and 1.2; we provide short proofs for completeness.